for some reason its saying I summited this already but I haven't hmmm
Find the standard deviation of the binomial random variable.

According to a college​ survey, 22% of all students work full time. Find the standard deviation for the random variable​ X, the number of students who work full time in samples of size 16. Round your answer to two decimal places
I used the square root *22*16*.84 but its not right am I using the right formula ?

Answer by addingup(3677)

The binomial random variable has an event that either occurs or does not occur. If it occurs we call it a success. If it doesn't we call it a failure.
In your problem P(full-time) is a success.
p=0.22
n = 16
Therefore, the expected number of successes:
np = 16(0.22) = 3.52
Variance = np(1-p) = 16(0.22)(1-0.22) = 16(0.22)(0.78) = 2.7456
Standard deviation: sqrt of variance = sqrt(2.7456) = 1.657